Market Top Investors, Inc., is considering the purchase of a $480,000 computer with an economic life of six years. The computer will be fully depreciated over six years using the straight-line method, at which time it will be worth $138,000. The computer will replace two office employees whose combined annual salaries are $99,000. The machine will also immediately lower the firm’s required net working capital by $88,000. This amount of net working capital will need to be replaced once the machine is sold. The corporate tax rate is 23 percent. The appropriate discount rate is 9 percent. Calculate the NPV of this project. (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.) Is it worthwhile to buy the computer? No Yes
Answer #1

Answer: NPV : $ 43,389.16

Yes, it is worthwhile to buy the computer.

Initial investment = $ 480,000 - $ 88,000 = $ 392,000.

Annual depreciation = $ 480,000 / 6 = $ 80,000

Earnings before Taxes, Depreciation and Amortization ( EBITDA) = 99,000.

Annual Operating Cash Flows After Taxes = EBITDA x ( 1 - T ) + Depreciation x T = 99,000 x 0.77 + 80,000 x 0.23 = 76,230 + 18,400 = 94,630.

After tax salvage value of computer = $ 138,000 x 0.77 = $ 106,260

Terminal cash flows = $ 106,260 - $ 88,000 = $ 18,260

PVA 9%, 6 years = [ { 1 - ( 1 / ( 1.09 ) 6 } / 0.09 ] = 4.4859

PV9%, 6th year = ( 1 / 1.09 ) 6 = 0.5963

NPV = 94,630 x 4.4859 + 18,260 x 0.5963 - 392,000 = 424,500.72 + 10,888.44 - 392,000 = 43,389.16